OpenStudy (anonymous):

If f(x)=3x+8, for what is the value of "a" is f(a)=a?

OpenStudy (anonymous):

you mean what is the value of a when f(a)=a

OpenStudy (anonymous):

\[f(a)=3(a)+8\] but \[f(a)=a\] so they are equall??????

OpenStudy (anonymous):

-4?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yes that would work

OpenStudy (anonymous):

@Jonask a=3a+8 -8=2a a=-4

OpenStudy (anonymous):

yes thats how we can work it out,cool